OpenStudy (annie96):

Question 1.1. Choose the expression that forms an equation with the given expression. 8 • 7 – 4 = (2 + 9) • 2 10 • 3 – 6 9 • 5 + 7 9(12 ÷ 2) Question 2.2. Choose the expression that completes the equation. 30 ÷ 2 + 10 = ________ (Points : 1) 3(2 + 3) 7(2 + 1) 6(3 + 7) 5(4 +1)

OpenStudy (texaschic101):

well....8 * 7 - 4 = 56 - 4 = 52 and 9 * 5 + 7 = 45 + 7 = 52 ===================== 30/2 + 10 = 15 + 10 = 25 and 5(4 + 1) = 5 * 5 = 25

OpenStudy (annie96):

Thank you so much:).

OpenStudy (texaschic101):

sure thing :)